.Determine how many bit strings of length 5 can be formed ,where three consecutive 0s are not allowed .


1
Expert's answer
2021-08-16T13:57:09-0400

The number of elements of the set of different binary strings of length 5 is "2^5=32."

Determine the number of bit strings of length 5 that contain three consecutive 0s.


"\begin{matrix}
 0 & 0 & 0 & 0 & 0\\
 0 & 0 & 0 & 0 & 1\\
 0 & 0 & 0 & 1 & 0\\
 0 & 0 & 0 & 1 & 1\\
\\
 1 & 0 & 0 & 0 & 0\\
 1 & 0 & 0 & 0 & 1\\
\\
 0 & 1 & 0 & 0 & 0\\
 1 & 1 & 0 & 0 & 0\\
\end{matrix}"

How many bit strings of length 5 can be formed ,where three consecutive 0s are not allowed


"32-8=24"